OpenStudy (anonymous):

could you show work (3x^2+2x-5) + (2x^2-8x+3)

OpenStudy (anonymous):

3x^2 +2x^2 = 5x^2 2x-8x = -6x -5+3= -2 ---->5x^2-6x-2